OpenStudy (anonymous):

if f(x)=x^-1/3 then find the value of f(1000)

OpenStudy (anonymous):

plug 1000 in for x

OpenStudy (anonymous):

the denominator is the root so the cube root of x and the negative sign makes it the reciprocal so \[\frac{ 1 }{ \sqrt[3]{1000} }=1/10\]

OpenStudy (anonymous):

:) you are so kind'

OpenStudy (anonymous):

no problem. I understand you are in a crisis so I am actually solving quicker than I normally would. But I am still trying to help you understand it

OpenStudy (anonymous):

yes u are doing amazing